15-n=-1-4(7n-4)
Solve for n

Found 2 solutions by rfer, ikleyn:
Answer by rfer(16322)

15-n=-1-28n+16
-n+28n=-1+16-15
27n=0
n=0

Answer by ikleyn(52781)

.
If you have difficulties solving so simple equations, read the lesson
    - HOW TO solve a linear equation
in this site.